# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ 8. Weak solutions of ISDE for the lower Dirichlet form
i The quasi-regularity of Dirichlet forms is a critical property for applying stochastic analysis to these objects. From Theorem 11.1 and the general theory of quasi-regular Dirichlet forms (cf. [5]), we have diffusions that are properly associated with Dirichlet forms in (11.6) and (11.7).
ii The proof of the sub-diffusivity of the tagged particles of the Ginibre interacting Brownian motion (the two-dimensional Coulomb interacting Brownian motion with $\beta=2$) relies on (11.6) [28, 29]. This result reflects the effect of the long-range nature of the two-dimensional Coulomb potential. Indeed, the tagged particles of interacting Brownian motion in $\mathbb{R}^{d}$ under Ruelle’s class potential with a convex hardcore are always diffusive if $d\geq 2$ [23].
iii Suzuki proved that the unlabeled Markov processes associated with a Dirichlet form are ergodic if the reference measure $\mu$ is number rigid and the Dirichlet form is locally ergodic [40]. This theorem can be applied to the Ginibre random point field. In [40], Suzuki considered the Dirichlet form that corresponds to $(\underline{\mathscr{E}}_{\star\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star\star}^{\mu^{[m]}})$. Hence, from Theorem 11.1, we see that the diffusion associated with the Dirichlet form for the Ginibre random point field is ergodic and the Markov process in [40] is a diffusion.
iv Suzuki [41] studied curvature bound of Dyson’s Brownian motion, which is an infinite particle system of interacting Brownian particles in $\mathbb{R}$ with the two-dimensional Coulomb potential with the inverse temperature $\beta=2$. Thus, the dynamics have a Riesz potential. He studied Dirichlet form of the type $(\underline{\mathscr{E}}_{\star\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star\star}^{\mu^{[m]}})$. Our argument for Theorem 11.1 is valid for the Riesz potential. Thus, his result is valid for the dynamics of Dyson’s model studied by [25, 34] because of the uniqueness of Dirichlet forms.
iv In [36, 30], the irreducibility and ergodicity of Dyson’s model were proved. The proof is also valid for the Coulomb interacting Brownian motions in the present article. .
For nonnegative, symmetric bilinear forms, we write $(\mathscr{E}_{1},\mathscr{D}_{1})\prec(\mathscr{E}_{2},\mathscr{D}_{2})$ if $(\mathscr{E}_{2},\mathscr{D}_{2})$ is an extension of $(\mathscr{E}_{1},\mathscr{D}_{1})$, that is, $\mathscr{D}_{1}\subset\mathscr{D}_{2}$ and $\mathscr{E}_{1}(f,g)=\mathscr{E}_{2}(f,g)$ for all $f,g\in\mathscr{D}_{1}$. Then, by construction, we can easily deduce that, for all $\mathsf{a}\in\mathsf{S}_{0}$,
$(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\overline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})\prec$ $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})\prec(\underline{\mathscr{E}}_{\star}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu_{\mathsf{a}}^{[m]}}),$
$(\mathscr{E}^{\mu^{[m]}}_{\star},\overline{\mathscr{D}}^{\mu^{[m]}}_{\star})\prec$ $(\underline{\mathscr{E}}_{\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star}^{\mu^{[m]}})\prec(\underline{\mathscr{E}}_{\star\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star\star}^{\mu^{[m]}}).$ (11.10)
Thus, we have proved in Theorem 11.1 that the relations in (11.10) are indeed equalities.
11.2. Proof of Theorem 2.4.
@@
Let $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ be as in Theorem 10.6. In the proof of Theorem 2.3, we constructed $\mathbf{X}$ as $\mathbf{X}:=\underline{\mathbf{X}}_{\mathsf{a}}=\overline{\mathbf{X}}_{\mathsf{a}}$ by proving $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ are indistinguishable. Let $\mathbf{X}^{[m]}$ be the $m$-labeled process of $\mathbf{X}$ defined by (2.42). In Theorem 2.3, we have already proved that $\mathbf{X}^{[m]}$ is a $\mu_{\mathsf{a}}^{[m]}$-symmetric diffusion for $\mu$-a.s. $\mathsf{a}$, but this does not necessarily mean the collection over $\mathsf{a}\in\mathsf{S}$ is a $\mu^{[m]}$-symmetric diffusion. Indeed, denoting by $\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]}$ the state space of the $\mu_{\mathsf{a}}^{[m]}$-symmetric diffusion $\mathbf{X}^{[m]}$, we have $\mu_{\mathsf{a}}^{[m]}((\mathbb{R}^{d})^{m}\times\mathsf{S}\backslash\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]})=0$ but it does not necessarily hold
$\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]}\cap\widetilde{\mathbf{S}}_{\mathsf{b}}^{[m]}=\emptyset\quad\text{ for }[\mathsf{a}]\neq[\mathsf{b}],$ (11.11)
where $\mathsf{a}]$ is the equivalent class in (2.51). If (11.11) holds, then the collection $\mathbf{X}^{[m]}$ over $\mathsf{a}$ is a diffusion with state space $\cup_{\mathsf{a}}\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]}$ because each $\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]}$ is the invariant set of the conservative diffusion $\mathbf{X}^{[m]}$. In the following theorem, we prove (11.11) and that the $m$-labeled process $\mathbf{X}^{[m]}$ is a diffusion with state space $(\mathbb{R}^{d})^{m}\times\mathsf{S}$.
Theorem 11.2.
Assume ((A1))((A3)). Let $\mathbf{X}$ and $\mathbf{X}^{[m]}$ be as above.
@@
From the proof of Theorem 2.3, we have already proved that $\mathbf{X}^{[m]}$ is a $\mu_{\mathsf{a}}^{[m]}$-symmetric, conservative diffusion for $\mu$-a.s. $\mathsf{a}$. From Theorem 10.6, $\mathbf{X}^{[m]}$ is properly associated with the Dirichlet form on $L^{2}(\mu^{[m]})$ in (11.6). Thus, we obtain i.
Let $\mathsf{K}_{\mathsf{a}}^{n}$ be the compact set given in the proof of Theorem 11.1. Let $\mathsf{a}]$ be as in (2.51). Recall that $\mathsf{K}_{\mathsf{a}}^{n}$ is increasing in $n$ for each $\mathsf{a}$ and that $\mathsf{K}_{\mathsf{a}}^{n}\subset[\mathsf{a}]$. Then,
$(\mathbb{R}^{d})^{m}\times\mathsf{K}_{\mathsf{a}}^{n}\subset(\mathbb{R}^{d})^{m}\times[\mathsf{a}]\quad\text{ for all }n\in\mathbb{N}.$ (11.12)
Let $\mathbf{K}_{\mathsf{a}}^{[m]}=(\mathbb{R}^{d})^{m}\times\cup_{n=1}^{\infty}\mathsf{K}_{\mathsf{a}}^{n}$ and $\mathbf{S}_{\mathsf{a}}^{[m]}=(\mathbb{R}^{d})^{m}\times[\mathsf{a}]$. From (11.9), $\mathbf{K}_{\mathsf{a}}^{[m]}=\mathbf{K}_{\mathsf{b}}^{[m]}$ if and only if $\mathsf{a}]=[\mathsf{b}]$. From (11.12), $\mathbf{K}_{\mathsf{a}}^{[m]}\cap\mathbf{K}_{\mathsf{b}}^{[m]}=\emptyset$ if $\mathsf{a}]\neq[\mathsf{b}]$. Indeed, this follows from $\mathsf{K}_{\mathsf{a}}^{n}\cap\mathsf{K}_{\mathsf{b}}^{o}=\emptyset$ for any $\mathsf{a}]\neq[\mathsf{b}]$ and all $n,o\in\mathbb{N}$, which follows from $\mathsf{K}_{\mathsf{a}}^{n}\subset[\mathsf{a}]$.
From the proof of Theorem 11.1, $\mathbf{K}_{\mathsf{a}}^{[m]}$ is the state space of the diffusion $\underline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ properly associated with $(\mathscr{E}^{\mu_{\mathsf{a}}^{[m]}},\underline{\mathscr{D}}^{\mu_{\mathsf{a}}^{[m]}})$ on $L^{2}(\mathbf{S}_{\mathsf{a}}^{[m]},\mu_{\mathsf{a}}^{[m]})$. Thus, $\mathbf{K}_{\mathsf{a}}^{[m]}$ is an invariant set of the diffusion $\underline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ for each $\mathsf{a}$. Because $\underline{\mathbf{X}}^{[m]}$ coincides with $\underline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ on $\mathbf{K}_{\mathsf{a}}^{[m]}$, $\underline{\mathbf{X}}^{[m]}$ is a $\mu^{[m]}$-symmetric, conservative diffusion properly associated with the superposition $(\underline{\mathscr{E}}_{\star\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star\star}^{\mu^{[m]}})$ on $L^{2}(\mu^{[m]})$ with the state space $\cup_{[\mathsf{a}]}\mathbf{K}_{\mathsf{a}}^{[m]}$ such that $\{\mathbf{K}_{\mathsf{a}}^{[m]}\}_{\mathsf{a}}$ gives a partition of the state space into $\mathbf{K}_{\mathsf{a}}^{[m]}$ that are invariant sets of $\underline{\mathbf{X}}^{[m]}$. We have thus proved ii. ∎
Proof of Theorem 2.4 .
Let $\mathbf{X}$ be the fully labeled process in Theorem 2.3. Let $\mathsf{X}=\mathbf{X}^{[0]}$ be the zero-labeled process given by $\mathbf{X}$. Recall that $\mu=\mu^{[0]}$. Then, from Theorem 11.2 ii, $\mathsf{X}$ is a $\mu$-reversible diffusion with state space $\mathsf{S}_{\star\star}$ properly associated with the Dirichlet form $(\underline{\mathscr{E}}_{\star\star}^{\mu},\underline{\mathscr{D}}_{\star\star}^{\mu})$ on $L^{2}(\mu)$. The state space $\mathsf{S}_{\star\star}$ is given by
@@
From Theorem 11.1 ii, $(\underline{\mathscr{E}}_{\star\star}^{\mu},\underline{\mathscr{D}}_{\star\star}^{\mu})$ on $L^{2}(\mu)$ is strongly local and quasi-regular. The state space $\mathsf{S}_{\star\star}$ satisfies $\mathrm{Cap}^{\mu}(\mathsf{S}_{\star\star}^{c})=0$, where $\mathrm{Cap}^{\mu}$ is the capacity associated with $(\underline{\mathscr{E}}_{\star\star}^{\mu^{[m]}},\underline{\mathscr{D}}_{\star\star}^{\mu^{[m]}})$ on $L^{2}(\mu)$. This follows from that $\mathbf{K}_{\mathsf{a}}^{[0]}$ is an invariant set of the diffusion $\underline{\mathsf{X}}_{\mathsf{a}}=\underline{\mathbf{X}}_{\mathsf{a}}^{[0]}$ for each $\mathsf{a}\in\mathsf{S}_{0}$. Because $\mathsf{X}=\underline{\mathsf{X}}_{\mathsf{a}}$, $\mathbf{K}_{\mathsf{a}}^{[0]}$ is the invariant set of the diffusion $\mathsf{X}$ for each $\mathsf{a}\in\mathsf{S}_{0}$.
@@
Let $\mathfrak{l}_{\mathrm{path}}\!:\!W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})\!\to\!C([0,\infty);\mathbb{R}^{d})^{\mathbb{N}}$ be as in Lemma 13.2. Let $\mathfrak{u}_{\mathrm{path}}$ be the map defined by $\mathfrak{u}_{\mathrm{path}}(\mathbf{w})=\mathsf{w}$, where $\mathsf{w}(t)=\sum_{i}\delta_{w^{i}(t)}$ for $\mathbf{w}=(w^{i})_{i}$. Then, $\mathfrak{u}_{\mathrm{path}}\circ\mathfrak{l}_{\mathrm{path}}$ is the identity map. Hence, $\mathfrak{l}_{\mathrm{path}}$ is injective. We note that $\mathbf{X}=\mathfrak{l}_{\mathrm{path}}(\mathsf{X})$.
@@
Let $P_{\mathbf{x}}=P(\mathbf{X}\in\cdot|\mathbf{X}_{0}=\mathbf{x})$ be the distribution of $\mathbf{X}$ starting at $\mathbf{x}\in\mathfrak{u}^{-1}(\mathsf{S}_{\star\star})$. As before, let $\mathsf{P}_{\mathsf{x}}$ be the distribution of $\mathsf{X}$ starting at $\mathsf{x}$. Then, we have
@@
Let $\sigma$ be an arbitrary $\{\mathscr{F}_{t}\}$-stopping time such that $\sigma<\infty$ a.s. Let $\tau=\sigma\circ\mathfrak{l}_{\mathrm{path}}$ and $\mathscr{G}_{t}=\mathfrak{l}_{\mathrm{path}}^{-1}(\mathscr{F}_{t})$. Then, we have $\mathscr{G}_{\tau}=\mathfrak{l}_{\mathrm{path}}^{-1}(\mathscr{F}_{\sigma})$. Recall that $\mathfrak{l}_{\mathrm{path}}$ is injective and $\mathsf{X}$ is a diffusion with state space $\mathsf{S}_{\star\star}$. Then, for each $\mathbf{x}\in\mathbf{S}_{\star\star}$ such that $\mathbf{x}=\mathbf{X}_{\sigma(\mathbf{X})}$,
@@
12. Dynamical finite-volume approximations: Proof of Theorem 2.5.
In this section, we complete the proof of Theorem 2.5. Let $\underline{\mathbf{X}}_{R}^{N_{\mathfrak{n}}}$ be as in Theorem 2.5.

### v4

@@ 11.1. Uniqueness of extensions of Dirichlet forms
Kawamoto proved the identity (11.11) for Dyson’s model [13]:
$(\mathscr{E}^{[m]},\mathscr{D}_{\mathrm{upr}}^{[m]})=(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}}^{[m]}).$ (11.11)
This identity, together with the uniqueness theorem for solutions of ISDEs in [34], implies that the tail $\sigma$-field of $\mu$ is invariant under the dynamics of the associated diffusion. Recently, Kawamoto has announced an extension of this identity to a broader class of RPFs (see [15]). It is therefore plausible that the same identity also holds for the Coulomb RPFs.
11.2. Proof of Theorem 2.4
@@
Let $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ be as in Theorem 10.1. In the proof of Theorem 2.3, we constructed $\mathbf{X}$ as $\mathbf{X}:=\underline{\mathbf{X}}_{\mathsf{a}}=\overline{\mathbf{X}}_{\mathsf{a}}$ by proving $\underline{\mathbf{X}}_{\mathsf{a}}$ and $\overline{\mathbf{X}}_{\mathsf{a}}$ are indistinguishable. Let $\mathbf{X}^{[m]}$ be the $m$-labeled process of $\mathbf{X}$, see (2.34).
In Theorem 2.3, we have already proved that $\mathbf{X}^{[m]}$ is a $\mu_{\mathsf{a}}^{[m]}$-symmetric diffusion for $\mu$-a.s. $\mathsf{a}$. However, this does not necessarily imply that the family $\mathbf{X}^{[m]}$ indexed by $\mathsf{a}\in\mathsf{S}$ is a $\mu^{[m]}$-symmetric diffusion.
Indeed, let $\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]}$ be an invariant set of the $\mu_{\mathsf{a}}^{[m]}$-symmetric diffusion $\mathbf{X}^{[m]}$. Then $\mu_{\mathsf{a}}^{[m]}((\mathbb{R}^{d})^{m}\times\mathsf{S}\backslash\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]})=0$. However, it does not necessarily hold that
$\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]}\cap\widetilde{\mathbf{S}}_{\mathsf{b}}^{[m]}=\emptyset\quad\text{whenever }[\mathsf{a}]\neq[\mathsf{b}].$ (11.12)
where $\mathsf{a}]$ is the equivalence class introduced in (2.37).
If (11.12) holds, then the family $\mathbf{X}^{[m]}$ indexed by $\mathsf{a}$ is a diffusion with state space $\cup_{\mathsf{a}}\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]}$, because each $\widetilde{\mathbf{S}}_{\mathsf{a}}^{[m]}$ is an invariant set of the conservative diffusion $\mathbf{X}^{[m]}$. In the following theorem, we prove (11.12) for $\mathbf{K}_{\mathsf{a}}^{[m]}$, defined in the proof, and show that the $m$-labeled process $\mathbf{X}^{[m]}$ is a diffusion with state space $(\mathbb{R}^{d})^{m}\times\mathsf{S}$.
Proposition 11.1.
@@
By Theorem 10.1 ii, $\mathbf{X}^{[m]}$ is a $\mu_{\mathsf{a}}^{[m]}$-symmetric diffusion properly associated with $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{lwr}}^{[m]})$ on $L^{2}(\mu_{\mathsf{a}}^{[m]})$. Hence by (11.7), we obtain i.
Let $\mathsf{K}_{\mathsf{a}}^{n}$ be the compact set given in the proof of Theorem 11.1. Let $\mathsf{a}]$ be as in (2.37). $\mathsf{K}_{\mathsf{a}}^{n}$ is increasing in $n$ for each $\mathsf{a}$ and that $\mathsf{K}_{\mathsf{a}}^{n}\subset[\mathsf{a}]$ and
$(\mathbb{R}^{d})^{m}\times\mathsf{K}_{\mathsf{a}}^{n}\subset(\mathbb{R}^{d})^{m}\times[\mathsf{a}],\quad n\in\mathbb{N}.$ (11.13)
Let $\mathbf{K}_{\mathsf{a}}^{[m]}=(\mathbb{R}^{d})^{m}\times\cup_{n=1}^{\infty}\mathsf{K}_{\mathsf{a}}^{n}$. By (11.10), $\mathbf{K}_{\mathsf{a}}^{[m]}=\mathbf{K}_{\mathsf{b}}^{[m]}$ if and only if $\mathsf{a}]=[\mathsf{b}]$. By (11.13), $\mathbf{K}_{\mathsf{a}}^{[m]}\cap\mathbf{K}_{\mathsf{b}}^{[m]}=\emptyset$ if $\mathsf{a}]\neq[\mathsf{b}]$.
By the proof of Theorem 11.1, $\mathbf{K}_{\mathsf{a}}^{[m]}$ is the state space of the diffusion properly associated with $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ on $L^{2}((\mathbb{R}^{d})^{m}\times[\mathsf{a}],\mu_{\mathsf{a}}^{[m]})$. By Theorem 10.1 ii, $\underline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ is properly associated with $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$. Thus, $\mathbf{K}_{\mathsf{a}}^{[m]}$ is an invariant set of the diffusion $\underline{\mathbf{X}}_{\mathsf{a}}^{[m]}$.
Since $\underline{\mathbf{X}}^{[m]}$ coincides with $\underline{\mathbf{X}}_{\mathsf{a}}^{[m]}$ on $\mathbf{K}_{\mathsf{a}}^{[m]}$, $\underline{\mathbf{X}}^{[m]}$ is a $\mu^{[m]}$-symmetric, conservative diffusion properly associated with the superposition $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{upr}}^{[m]})$ of $(\mathscr{E}_{\mathsf{a}}^{[m]},\mathscr{D}_{\mathsf{a},\mathrm{upr}}^{[m]})$ on $L^{2}(\mu^{[m]})$ with the state space $\sqcup_{[\mathsf{a}]}\mathbf{K}_{\mathsf{a}}^{[m]}$. Here $\{\mathbf{K}_{\mathsf{a}}^{[m]}\}_{[\mathsf{a}]}$ gives a partition of the state space into $\mathbf{K}_{\mathsf{a}}^{[m]}$. Each $\mathbf{K}_{\mathsf{a}}^{[m]}$ is an invariant sets of $\underline{\mathbf{X}}^{[m]}$. We have thus proved ii. ∎
Proof of Theorem 2.4. Redefining $\mathsf{S}_{0}$ in Proposition 11.1 if necessary, we may assume that its conclusion holds for all $\mathsf{a}\in\mathsf{S}_{0}$, with $\mu(\mathsf{S}_{0})=1$.
Let $\mathbf{X}$ be the fully labeled process in Theorem 2.3. Let $\mathsf{X}=\mathbf{X}^{[0]}$ be the zero-labeled process given by $\mathbf{X}$. Recall that $\mu=\mu^{[0]}$. From Proposition 11.1 ii, $\mathsf{X}$ is a $\mu$-reversible diffusion with state space $\mathsf{S}_{\star\star}$ properly associated with the Dirichlet form $(\mathscr{E}^{\mu},\mathscr{D}_{\star,\mathrm{lwr}\star}^{\mu})$ on $L^{2}(\mu)$. $\mathsf{S}_{\star\star}$ is given by
@@
From Theorem 11.1 ii, $(\mathscr{E}^{\mu},\mathscr{D}_{\star,\mathrm{lwr}\star}^{\mu})$ on $L^{2}(\mu)$ is strongly local and quasi-regular. The state space $\mathsf{S}_{\star\star}$ satisfies $\mathrm{Cap}(\mathsf{S}_{\star\star}^{c})=0$, where $\mathrm{Cap}$ is the capacity associated with $(\mathscr{E}^{[m]},\mathscr{D}_{\star,\mathrm{lwr}\star}^{[m]})$ on $L^{2}(\mu)$. This follows from that $\mathbf{K}_{\mathsf{a}}^{[0]}$ is an invariant set of the diffusion $\underline{\mathsf{X}}_{\mathsf{a}}=\underline{\mathbf{X}}_{\mathsf{a}}^{[0]}$ for each $\mathsf{a}\in\mathsf{S}_{0}$. Because $\mathsf{X}=\underline{\mathsf{X}}_{\mathsf{a}}$, $\mathbf{K}_{\mathsf{a}}^{[0]}$ is the invariant set of the diffusion $\mathsf{X}$ for each $\mathsf{a}\in\mathsf{S}_{0}$.
@@
Let $\mathfrak{l}_{\mathrm{path}}\!:\!W_{\mathrm{NE}}(\mathsf{S}_{\mathrm{s,i}})\!\to\!W(\mathbb{R}^{d})^{\mathbb{N}}$ be as in Lemma 13.2. Let $\mathfrak{u}_{\mathrm{path}}$ be the map defined by $\mathfrak{u}_{\mathrm{path}}(\mathbf{w})=\mathsf{w}$, where $\mathsf{w}(t)=\sum_{i}\delta_{w^{i}(t)}$ for $\mathbf{w}=(w^{i})_{i}$. $\mathfrak{u}_{\mathrm{path}}\circ\mathfrak{l}_{\mathrm{path}}$ is the identity map. Hence, $\mathfrak{l}_{\mathrm{path}}$ is injective. We note that $\mathbf{X}=\mathfrak{l}_{\mathrm{path}}(\mathsf{X})$.
@@
Let $P_{\mathbf{x}}=P(\mathbf{X}\in\cdot|\mathbf{X}_{0}=\mathbf{x})$ be the distribution of $\mathbf{X}$ starting at $\mathbf{x}\in\mathfrak{u}^{-1}(\mathsf{S}_{\star\star})$. As before, let $\mathsf{P}_{\mathsf{x}}$ be the distribution of $\mathsf{X}$ starting at $\mathsf{x}$. We have
@@
Let $\sigma$ be an arbitrary $\{\mathscr{F}_{t}\}$-stopping time such that $\sigma<\infty$ a.s. Let $\tau=\sigma\circ\mathfrak{l}_{\mathrm{path}}$ and $\mathscr{G}_{t}=\mathfrak{l}_{\mathrm{path}}^{-1}(\mathscr{F}_{t})$. We have $\mathscr{G}_{\tau}=\mathfrak{l}_{\mathrm{path}}^{-1}(\mathscr{F}_{\sigma})$. $\mathfrak{l}_{\mathrm{path}}$ is injective and $\mathsf{X}$ is a diffusion with state space $\mathsf{S}_{\star\star}$. For each $\mathbf{x}\in\mathbf{S}_{\star\star}$ such that $\mathbf{x}=\mathbf{X}_{\sigma(\mathbf{X})}$,
@@
12. Finite-volume approximation: Proof of Theorem 2.5
All results in Section 12 are proved under Assumptions ((A1))((A3)) and (2.44)–(2.45).
12.1. Finite-domain approximation
We now impose reflecting boundary conditions on $\partial S_{R}$ in (2.42) and fix the particles outside the region $\overline{S}_{R}=\{s\in\mathbb{R}^{d};|s|\leq R\}$. We suppose $S_{R}\subset O_{N}$. Then the dynamics are described as follows:
$\underline{X}_{R}^{N,i}(t)-\underline{X}_{R}^{N,i}(0)$ $=\int_{0}^{t}\sigma(\underline{X}_{R}^{N,i}(u))\,dB_{u}^{i}+\int_{0}^{t}\mathfrak{b}^{N}\Big(\underline{X}_{R}^{N,i}(u),\sum_{j\neq i}^{N}\delta_{\underline{X}_{R}^{N,j}(u)}\Big)du$
$\quad+\int_{0}^{t}\frac{1}{2}\mathbf{n}_{R}\mathfrak{a}(\underline{X}_{R}^{N,i}(u))\,dL_{R}^{N,i}(u),$
$L_{R}^{N,i}(t)$ $=\int_{0}^{t}1_{\partial S_{R}}(\underline{X}_{R}^{N,i}(u))\,dL_{R}^{N,i}(u),\quad i\leq\mathsf{s}(\overline{S}_{R}),$ (12.1)
$\underline{X}_{R}^{N,i}(t)-\underline{X}_{R}^{N,i}(0)$ $=0,\quad\mathsf{s}(\overline{S}_{R})<i.$ (12.2)
Here $\frac{1}{2}\mathbf{n}_{R}\mathfrak{a}$ denotes the inward unit normal vector on $\partial S_{R}$, and $L_{R}^{N,i}$ is the local time on the boundary $\partial S_{R}$.
Let $\{N_{\mathfrak{n}}\}_{n=1}^{\infty}$ be the sequence in ((A2)) such that $\mu^{N_{\mathfrak{n}}}$ converges to $\mu$. Taking the limit $N_{\mathfrak{n}}\to\infty$ in the SDE (12.1)–(12.2), we obtain the ISDE (8.1)–(8.2).
Remark 12.1.
i The solutions of the above SDE remains fixed outside the region $\overline{S}_{R}$. This follows from (2.22) and (12.2).
ii Let $\mathbf{s}=(s^{i})_{i=1}^{N}$ satisfy $s^{i}\neq s^{j}$ for all $i\neq j$. Then the SDE (12.1)–(12.2) with $\underline{\mathbf{X}}_{R}^{N}(0)=\mathbf{s}$ admits a unique strong solution $\underline{\mathbf{X}}_{R}^{N}$. Indeed, the coefficients are smooth outside the collision set $\mathscr{N}=\bigcup_{i\neq j}\{s^{i}=s^{j}\}$, and $\underline{\mathbf{X}}_{R}^{N}$ does not hit $\mathscr{N}$ since $d\geq 2$; see Proposition 6.1.